OpenStudy (anonymous):

Sal climbs stairs by taking either one, two, or three steps at a time. Determine a recursive formula for the number of different ways Sal can climb a flight of n steps. Don't forget to include the initial conditions

OpenStudy (anonymous):

wow who knows' lets try it 1 step, 1 way 2 steps, 2 ways 3 steps, 4 ways (i think)

OpenStudy (anonymous):

how about 5 steps?

OpenStudy (anonymous):

oh lets try 4

OpenStudy (anonymous):

How did you get four ways for three steps? I got 3 one-steps, 1 two-step + 1 one-step, or 1 three-step?

OpenStudy (anonymous):

for three steps i though 1 1 1 1 2 2 1 3

OpenStudy (anonymous):

maybe this has something to do with partitions, i don't know the answer, just thinking out loud

OpenStudy (anonymous):

Oh that's a very good point! I didn't think about switching the pattern.

OpenStudy (anonymous):

4 steps 1 1 1 1 1 1 2 1 2 1 2 1 1 1 3 3 1

OpenStudy (anonymous):

see what you can do, i will try in a minute, but unless this is some application of specific topic, i think experimentation is the way to go

OpenStudy (anonymous):

5 steps 11111 1112 1121 1211 2111 113 131 311 Right?

OpenStudy (anonymous):

If we're going with experimentation...

OpenStudy (anonymous):

Oh or there's 221 to add to that taking five steps part

OpenStudy (anonymous):

And for four stairs, you can also take 2 2, so so far, we have a pattern of 1, 2, 4, 7, 9

OpenStudy (anonymous):

i think your five steps is missing something

OpenStudy (anonymous):

1 2 2 2 1 2 1 2 2 2 3 3 2

OpenStudy (anonymous):

So far I have: 11111 1112 1121 1211 2111 113 131 311 221 122

OpenStudy (anonymous):

So 12 for taking five steps?

OpenStudy (anonymous):

1 1 1 1 1 1 1 1 2 1 1 2 1 1 2 1 1 2 1 1 1 1 1 3 1 3 1 3 1 1 2 3 3 2

OpenStudy (anonymous):

oh right i missed a couple also 2 2 1 2 1 2 1 2 2

OpenStudy (anonymous):

So now our pattern so far is 1, 2, 4, 7, 12...

OpenStudy (anonymous):

lets see what we have so far 1 gives 1 2 gives 2 3 gives 4 4 gives 7 5 gives 13

OpenStudy (anonymous):

Can you write out what pattern you got for 5? I only have 12 for some reason...must be forgetting one?

OpenStudy (anonymous):

1 1 1 1 1 1 1 1 2 1 1 2 1 1 2 1 1 2 1 1 1 1 2 2 2 1 2 2 2 1 1 1 3 1 3 1 3 1 1 3 2 2 3

OpenStudy (sirm3d):

let's say there are x 3-step-at-a-time, y 2-step-at-time, and z one-step.

OpenStudy (anonymous):

yes it is time to cut to the chase, i am sure there is a snappier way to do this

OpenStudy (sirm3d):

we can permute the 3-step, 2-step, and 1-step into (x+y+z)!/z!y!z!) ways

OpenStudy (anonymous):

oh and also we are looking for a recursion , not a direct formula in terms of \(n\) although that might be helpful

OpenStudy (sirm3d):

oh, yeah. i forgot that part. i was into counting the ways.

OpenStudy (anonymous):

Yes...the requirement is that we find a recursive formula for the number of different ways Sal can climb a flight of n steps.

OpenStudy (anonymous):

Do we need an initial condition?

OpenStudy (anonymous):

in other words, when we add one more step, can we figure out how many more ways we need

OpenStudy (anonymous):

So you want to figure out how many ways for six stairs?

OpenStudy (anonymous):

maybe it is just \(a_n=2a_{n-1}-1\) wouldn't that be easy ?

OpenStudy (anonymous):

i got 13 for 5 steps lets see if we get 25 for 6 steps

OpenStudy (anonymous):

Ha that would be crazy if the definition would be that easy :)

OpenStudy (anonymous):

Let's see...for six steps I get: 111111 11112 11121 11211 12111 21111 1122 1221 2211 1212 2121 33 1113 1131 1311 3111

OpenStudy (anonymous):

Is this right?

OpenStudy (anonymous):

I think that sirm3D had a good idea by assigning the values to one-step, two-step, and three-step

OpenStudy (anonymous):

missing 1 2 3 1 3 2 2 1 3 2 3 1 3 1 2 3 2 1

OpenStudy (anonymous):

Oh good catch! So we're at 22 now

OpenStudy (anonymous):

i think this may be a very hard problem without a gimmick maybe we can see what happens to the list when you add 1

OpenStudy (sirm3d):

consider n = 6: (111111) = 1, (11112) = 5, (1122) = 6, (1113) = 4, (123) = 6, (222) = 1, (33)=1. Total of 24 possible ways

OpenStudy (anonymous):

also missing 2 2 2

OpenStudy (sirm3d):

i did not. there's (222)

OpenStudy (anonymous):

sorry i meant earlier

OpenStudy (anonymous):

sirm3d what are your six ways of using 1122?

OpenStudy (anonymous):

I got 1122 1221 2211 1212 2121 Which one am I missing?

OpenStudy (anonymous):

think of it as \(\dbinom{4}{2}=6\)

OpenStudy (anonymous):

missing 2 1 1 2

OpenStudy (sirm3d):

4!/(2!2!)

OpenStudy (anonymous):

so much for my \(2a_{n-1}-1\) theory

OpenStudy (anonymous):

Oh yeah..that's a very good point. Well the theory you had, satellite73 was definitely an easier one than I think we're faced with right now :/

OpenStudy (anonymous):

maybe one more n = 7 (1111111) =1 (111112) = 7 (11122) = 10 (1 2 2 2) = 4 (1 1 1 1 3) =5 (2 2 3) = 3 (1 1 2 3) = 8 i have no damned idea

OpenStudy (sirm3d):

(111112) = 6 only.

OpenStudy (anonymous):

I think rather than trying to do this out by brute force, we should look at the pattern and see if a formula can't be found...?

OpenStudy (sirm3d):

i found it. ^^

OpenStudy (sirm3d):

\[\large a_1=1,\quad a_2=2, \quad a_3=4\]

OpenStudy (anonymous):

Oh good! So those are your conditions..

OpenStudy (sirm3d):

@rachelk09 @satellite73 \[\large a_n=a_{n-1}+a_{n-2}+a_{n-3}, \quad n=4,5,6,...\]

OpenStudy (anonymous):

Let's try using that to see if it works for the six that we did out long hand...looks right to me...

OpenStudy (anonymous):

The first one was 1 + 2 + 4, which does equal seven. Atempting the next set now. I think you got it!!

OpenStudy (sirm3d):

n=1: 1 n=2: (11) = 1, (2) = 1 n=3: (111)=1, (12)=2, (3) = 1 n=4: (1111)=1, (112)=3, (13)=2, (22)=1 n=5: (11111)=1, (1112)=4, (122)=3, (113)=3, (23)=2 n=6: (111111)=1, (11112)=5, (1122)=6, (1113)=4, (123)=6, (222)=1, (33)=1 n=7: (1111111)=1, (111112)=6, (11122)=10, (1222)=4, (11113)=5, (1123)=12, (133)=3, (223)=3

OpenStudy (anonymous):

Everything checks out. You did it! Thanks so so much!

OpenStudy (sirm3d):

YW. ^^